Prepare the entry, if necessary, to adjust the account balances because of the revision of the estimated life in 2019. (If noExercise 11-12 In 1990, Whispering Company completed the construction of a building at a cost of $2,480,000 and first occupied it in January 1991. It was estimated that the building will have a useful life of 40 years and a salvage value of $74,400 at the end of that time. Early in 2001, an addition to the building was constructed at a cost of $620,000. At that time, it was estimated that the remaining life of the building would be, as originally estimated, an additional 30 years, and that the addition would have a life of 30 years and a salvage value of $24,800. In 2019, it is determined that the probable life of the building and addition will extend to the end of 2050, or 20 years beyond the original estimate. Using the straight-line method, compute the annual depreciation that would have been charged from 1991 through 2000. Annual depreciation from 1991 through 2000 $ / yr. SHOW LIST OF ACCOUNTS LINK TO TEXT LINK TO TEXT Compute the annual depreciation that would have been charged from 2001 through 2018. Annual depreciation from 2001 through 2018 $ / yr. SHOW LIST OF ACCOUNTS LINK TO TEXT LINK TO TEXT Prepare the entry, if necessary, to adjust the account balances because of the revision of the estimated life in 2019. (If no entry is required, select "No entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.) Account Titles and Explanation Debit Credit SHOW LIST OF ACCOUNTS LINK TO TEXT LINK TO TEXT Compute the annual depreciation to be charged, beginning with 2019. (Round answer to 0 decimal places, e.g. 45,892.) Annual depreciation expense—building $ Click if you would like to Show Work for this question: Open Show Work
Answer #1
1
Cost 2480000
Less: Salvage value 74400
Depreciable cost 2405600
Useful life 40
Annual depreciation from 1991 through 2000 60140 /yr.
2
Cost 2480000
Add: Additions 620000
Total cost 3100000
Less: Accumulated depreciation from 1991 through 2000 601400 =60140*10
Book value, Jan 2001 2498600
Less: Salvage value 99200 =74400+24800
Depreciable cost 2399400
Useful life 30
Annual depreciation from 2001 through 2018 79980 /yr.
3
No entry is required because of the revision of the estimated life in 2019
No entry                       0
         No entry                            0
4
Book value, Jan 2001 2498600
Less: Accumulated depreciation from 2001 through 2018 1439640 =79980*18
Book value, Jan 2019 1058960
Less: Salvage value 99200 =74400+24800
Depreciable cost 959760
Useful life 32 =50-18
Annual depreciation expense—building 29993